# TACL: Threshold-Adaptive Curriculum Learning Strategy for Enhancing Medical Text Understanding

### v1

@@ Abstract
Medical texts, particularly electronic medical records (EMRs), are a cornerstone of modern healthcare, capturing critical information about patient care, diagnoses, and treatments. These texts hold immense potential for advancing clinical decision-making and healthcare analytics. However, their unstructured nature, domain-specific language, and variability across contexts make automated understanding an intricate challenge. Despite the advancements in natural language processing, existing methods often treat all data as equally challenging, ignoring the inherent differences in complexity across clinical records. This oversight limits the ability of models to effectively generalize and perform well on rare or complex cases. In this paper, we present TACL (Threshold-Adaptive Curriculum Learning), a novel framework designed to address these challenges by rethinking how models interact with medical texts during training. Inspired by the principle of progressive learning, TACL dynamically adjusts the training process based on the complexity of individual samples. By categorizing data into difficulty levels and prioritizing simpler cases early in training, the model builds a strong foundation before tackling more complex records. By applying TACL to multilingual medical data, including English and Chinese clinical records, we observe significant improvements across diverse clinical tasks, including automatic ICD coding, readmission prediction and TCM syndrome differentiation. TACL not only enhances the performance of automated systems but also demonstrates the potential to unify approaches across disparate medical domains, paving the way for more accurate, scalable, and globally applicable medical text understanding solutions.
@@ I Introduction
Medical texts, such as electronic medical records (EMRs), are vital for modern healthcare, capturing patient care, diagnoses, and treatments [1]. However, their unstructured and specialized nature poses significant challenges for automated understanding. EMRs often contain domain-specific terminology, inconsistencies, and varying complexity—from straightforward cases to intricate ones involving rare conditions or multiple comorbidities.
The complexity of EMRs spans multiple dimensions, as depicted in Figure 1(a). Narrative structure ranges from explicit and coherent records to fragmented and ambiguous texts requiring deeper reasoning. Clinical variability introduces further challenges, as common cases are easier to address, whereas rare or complex cases demand advanced understanding. Additionally, patient demographics contribute to variability, such as age-related differences in medical documentation. Traditional NLP models often often adopt a ”one-size-fits-all” paradigm, treating all medical texts as equally challenging [2, 3, 4, 5, 6, 7], leading to suboptimal performance, particularly on rare and complex cases that hold significant clinical importance.
This raises a key question: can we design a training paradigm that allows models to gradually and systematically acquire knowledge, just as humans learn? Recently, curriculum learning [8], a training scheme that aims to let the model learn from easier to more complex samples or tasks, can benefit machine learning algorithms. The efficacy of replacing conventional training with curriculum learning has been demonstrated in many application fields of healthcare, such as medical image segmentation [9, 10] and automatic ICD coding [5]. Inspired by this, we innovatively function medical text understanding as a progressive task with curriculum learning based on text complexity. However, two critical challenges remain: 1) how to design an effective difficulty assessment process for medical texts, and 2) how to develop a curriculum learning strategy that progressively guides the model from easier to harder examples. Intuitively, we can define the difficulty of medical text using some simple metrics like sequence length or code frequency. However, these strategies fail to improve performance and may even degrade it by overlooking semantic nuances, as shown in Figure 1(b). Therefore, how to determine a reliable and interpretable way to quantify the complexity of clinical narratives remains an urgent challenge. Additionally, how to design adaptive training strategies based on different tasks instead of manually setting thresholds, in order to ensure both task-specific performance improvements and the ability to generalize across diverse medical datasets remains an issue.
@@
To address the two issues, we introduce TACL (Threshold-Adaptive Curriculum Learning) for medical text understanding, which addresses the challenges of quantifying text complexity and improving generalization across diverse medical datasets. To implement the curriculum strategy, we design a novel difficulty assessment method based on a domain-specific pre-trained language model, which effectively quantifies the complexity of clinical narratives by considering semantic, syntactic, and contextual features. Furthermore, we develop an adaptive training strategy that dynamically updates the curriculum to enhance the model’s ability to generalize across heterogeneous medical datasets.
@@
TACL is validated on multilingual, multi-domain datasets, including English (MIMIC-III [11], MIMIC-IV [12]) and Chinese clinical records, demonstrating state-of-the-art performance in tasks like readmission prediction, out-hospital mortality, ICD coding, and TCM syndrome differentiation [13]. TACL achieves robustness in handling rare and complex cases, advancing the state of medical text understanding.
The main contributions can be summarized as follows:
  1. 1.
@@
    To the best of our knowledge, this is the first work to apply curriculum learning to the medical text understanding task. By introducing curriculum learning into this domain, our approach provides a novel perspective for handling the inherent complexity and variability of clinical narratives.
  2. 2.
    We design a difficulty assessment method based on a domain-specific pre-trained language model, which effectively quantifies the complexity of medical texts. This method enables a systematic division of clinical narratives into difficulty levels, providing a structured foundation for the curriculum learning process.
  3. 3.
@@
    We propose an adaptive training strategy that dynamically adjusts the learning process based on the model’s performance. This strategy enhances the model’s ability to generalize across diverse medical datasets, reducing overfitting to specific domains and improving robustness in real-world applications.
@@ II Related Work
The challenges in automated medical text understanding have driven significant advancements in natural language processing (NLP) tailored for healthcare. This section reviews relevant work in two primary areas: curriculum learning and medical text understanding.
@@ II-A Curriculum Learning in NLP
Curriculum learning, inspired by the human learning process, structures the training process by exposing models to simpler examples before progressively introducing more complex ones. Bengio et al.[8] first proposed this paradigm, which has shown success in various NLP tasks, including text classification, machine translation, and question answering. For example, Platanios et al.[14] introduced a dynamic curriculum framework for neural machine translation, where difficulty levels were determined by the model’s confidence.
While these methods have demonstrated efficacy, most rely on static or predefined difficulty measures, limiting their adaptability to the inherent variability of real-world data. Recent works on adaptive curriculum learning [15, 16] dynamically assess sample difficulty during training, allowing models to adjust to changing data distributions. However, these approaches primarily target general NLP tasks and fail to address the unique challenges of domain-specific data, such as medical texts. To bridge this gap, TACL introduces a threshold-adaptive curriculum mechanism tailored to the complexity of medical data. By dynamically prioritizing samples based on their difficulty, TACL improves both task-specific performance and generalization across diverse datasets.
@@ II-B Medical Text Understanding
Automated medical text understanding plays a critical role in improving clinical decision-making and healthcare analytics. Early approaches relied on rule-based systems and traditional machine learning models, such as conditional random fields and support vector machines, which required extensive domain expertise and manual feature engineering. The rise of deep learning brought significant advancements, with CNNs and RNNs excelling in tasks like clinical text classification and named entity recognition.
Transformer-based models, such as BERT [17] and its domain-specific adaptations like ClinicalBERT [18] and BioBERT [19], have further advanced medical NLP by leveraging large-scale pretraining on biomedical corpora to capture domain-specific semantics. While effective, these models often treat all training samples equally, limiting their ability to handle rare or complex cases. TACL addresses this limitation by introducing a curriculum-based training strategy, enabling models to progressively learn from simpler to more complex cases and improving generalization.
Datasets like MIMIC-III [11] and MIMIC-IV [12] have been instrumental in advancing medical NLP, supporting tasks such as ICD coding and clinical risk prediction. However, the variability and complexity of these datasets highlight the need for adaptive frameworks like TACL to effectively handle heterogeneous and challenging data distributions.
@@ III-A Contextual Representation
  • English Tasks: For tasks such as ICD coding out-of-hospital mortality, and readmission prediction, we employ RoBERTa [20] pre-trained on medical corpora.
  • Chinese Tasks: For TCM syndrome differentiation, we utilize ZY-BERT [13] pre-trained on traditional Chinese medicine datasets.
@@ III-C Threshold-Adaptive Curriculum Learning (TACL)
The complete TACL framework is summarized in Algorithm 1.
Algorithm 1 Threshold-Adaptive Curriculum Learning
1:Input: Training data divided into difficulty levels $\mathcal{D}_{\mathrm{easy}},\mathcal{D}_{\mathrm{middle}},\mathcal{D}_{\mathrm{hard}}$, total epochs $E$, sliding window size NN, threshold coefficient $\beta$
2:Initialize: Current dataset $\mathcal{D}_{\mathrm{train}}=\mathcal{D}_{\mathrm{easy}}$, macro-F1 deque F=[]F=[\,] of length NN
3:for $t=1$ to $E$ do
4:  Train model on $\mathcal{D}_{\mathrm{train}}$ for one epoch
5:  Compute macro-F1 score $F_{\mathrm{current}}$
6:  Append $F_{\mathrm{current}}$ to $F$; discard oldest value if $|F|>N$
7:  if $|F|=N$ then
8:    Compute $\bar{\gamma}$ using equation (7)
9:    Compute $\gamma_{\Delta}$ using equation(8)
10:    if $\gamma_{\Delta}<\beta\cdot\bar{\gamma}$ then $\triangleright$ Check saturation
11:     if $\mathcal{D}_{\mathrm{train}}=\mathcal{D}_{\mathrm{easy}}$ then
12:      $\mathcal{D}_{\mathrm{train}}\leftarrow\mathcal{D}_{\mathrm{easy}}\cup\mathcal{D}_{\mathrm{middle}}$
13:     else if $\mathcal{D}_{\mathrm{train}}=\mathcal{D}_{\mathrm{easy}}\cup\mathcal{D}_{\mathrm{middle}}$ then
14:      $\mathcal{D}_{\mathrm{train}}\leftarrow\mathcal{D}_{\mathrm{easy}}\cup\mathcal{D}_{\mathrm{middle}}\cup\mathcal{D}_{\mathrm{hard}}$
15:     else
16:      Break: Early stop training if at hard stage
17:     end if
18:    end if
19:  end if
20:end for
21:Output: Trained model
@@ IV-A Datasets
We evaluate our framework across four tasks using publicly available benchmark datasets, with each dataset partitioned into three difficulty levels (Easy, Medium, and Hard) described in Section 3.2. Table I provides an overview of dataset statistics for all tasks.
Task 1: Patient Readmission: We predict whether a patient will be readmitted to the hospital within a specific time frame based on clinical notes. Using the ADMISSIONS and NOTEEVENTS tables from the MIMIC-III dataset [11], two subsets are constructed:
  • Discharge Summary: Each sample consists of detailed discharge notes from NOTEEVENTS table, making it a rich source of information for prediction tasks.
  • Early Notes: Besides of discharge summary, it also provides a snapshot of a patient’s condition 3 days earlier in their care, offering complementary insights compared to discharge summaries.
Task 2: Out-of-hospital Mortality: Using discharge summaries from the MIMIC-IV dataset [12], this binary classification task predicts whether a patient dies within 30 days after discharge, excluding in-hospital deaths.
Task 3: ICD Coding: We evaluate our model on three datasets for ICD coding tasks, with each dataset representing either ICD-9 or ICD-10 coding:
  • MIMIC-III Full (ICD-9): Contains 52,723 clinical documents annotated with 8,929 unique ICD-9 codes. Each document has a median of 14 codes, making it a highly multi-label classification problem.
  • MIMIC-III 50 (ICD-9): Focuses on the top 50 most frequent ICD-9 codes, including 11,368 clinical documents. Each document is annotated with a median of 5 codes, representing a simpler but well-established benchmark.
  • MIMIC-IV (ICD-10): Provides clinical notes annotated with more granular ICD-10 codes, making this task more challenging due to the increased code specificity.
Preprocessing steps for all datasets include converting text to lowercase, removing non-alphanumeric characters, and truncating to a maximum of 2,500 or 4,000 tokens, depending on the model’s input size constraints.
Task 4: TCM Syndrome Differentiation: We classify Traditional Chinese Medicine (TCM) syndromes based on descriptions of patient conditions using the TCM-SD dataset [13]. Unlike the English-based tasks above, this dataset is in Chinese, with its unique linguistic structure and domain-specific terminology adding significant complexity.
TABLE II: Hyperparameter used for the tasks.
Task Model Hyperparameters
Batch Size Learning Rate Optimizer Epochs Length $\beta$
Task 1
Task 2
Bi-LSTM 64 0.0001 Adam 30 1024 0.8
Bio-BERT 64 0.00005 BertAdam 20 512 0.6
ClinicalBERT 64 0.00005 BertAdam 20 512 0.7
Task 3 Bi-GRU 8 0.001 AdamW 20 4000 0.7
CNN 8 0.001 AdamW 20 4000 0.7
CAML 8 0.005 AdamW 60 4000 0.5
MultiResCNN 16 0.0005 AdamW 40 4000 0.9
LAAT 8 0.001 AdamW 30 4000 0.7
PLM-ICD 8 0.00005 AdamW 20 4000 0.7
Task 4 RoBERTa 128 0.00005 AdamW 30 512 0.7
TCM-BERT 128 0.00005 AdamW 30 512 0.7
ZY-BERT 128 0.00005 AdamW 30 512 0.8
@@ IV-B Implementation
Experiments are conducted using NVIDIA L40 48GB GPU. To ensure reproducibility, random seeds are fixed across all runs. The thresholds for growth rate in TACL are set from 0.5 to 0.8 according to the performance on validation subset, but may vary depending on the model. All relevant codes will be made publicly available [^1] https://anonymous.4open.science/r/medical-codes-354C.
For evaluating the performance of our proposed model, we employ three commonly used metrics: F1-score (Micro and Macro), AUC (Micro and Macro), and precision at K (P@K). The metrics are computed on the test set using the model that achieves the best performance on the validation set.
@@ IV-C Baseline
We compare the proposed method against a range of strong baselines across all tasks. For tasks involving clinical and biomedical text, such as readmission prediction, out-of-hospital mortality, and ICD coding, we evaluate models like Bi-LSTM [21], which utilizes pre-trained Word2Vec embeddings and bidirectional LSTM layers, and transformer-based models like BioBERT [22] and ClinicalBERT [18], both pre-trained on large biomedical corpora. For ICD coding, additional specialized models like CAML [23] with label attention mechanisms and MultiResCNN [24] for multi-scale feature extraction are employed, alongside PLM-ICD [25], which integrates a pre-trained BERT encoder with a label attention mechanism. For TCM syndrome differentiation, we evaluate Chinese-language models like RoBERTa-wwm [26], pre-trained with Whole Word Masking on large-scale Chinese text, and ZY-BERT [13], a TCM-specific model trained on a large unlabelled corpus, achieving state-of-the-art performance on TCM tasks.
@@ V-A Main Results
The performance of TACL is summarized across all four tasks in Figure 3, showcasing its consistent improvements over traditional training methods across various models and datasets. TACL enhances performance metrics such as Macro-F1, AUROC, and RP80, demonstrating its robustness and effectiveness in handling diverse medical text understanding challenges.
Notably, TACL provides significant gains for simpler models like Bi-GRU, CNN, and Bi-LSTM, particularly for tasks like readmission prediction and ICD coding, where these models often struggle with data complexity and rare labels. For example, TACL improves Macro-F1 scores and recall on rare codes, highlighting its ability to compensate for the limitations of simpler architectures.
TABLE III: Performance on Patient Readmission Prediction.
@@
Advanced models such as PLM-ICD, MultiResCNN, ClinicalBERT, and ZY-BERT also benefit from TACL, albeit with smaller but consistent gains. These improvements are especially pronounced in scenarios involving rare labels and imbalanced datasets, such as ICD-10 coding and TCM syndrome differentiation. TACL complements domain-specific pretraining by dynamically managing data complexity during training, leading to better generalization.
Overall, TACL’s consistent performance gains across both simpler and state-of-the-art models validate its adaptability and scalability for diverse tasks, including multilingual and domain-specific challenges. The results underscore its potential to enhance clinical decision-making and medical text understanding.
@@ V-B Analysis on Difficulty Definition
The effectiveness of curriculum learning depends heavily on how data difficulty is defined. Figure 1(b) and Table III show that traditional statistical features including sequence length, number of code, and frequency of code, degrade performance compared to the baseline. These results indicate that statistical features fail to effectively capture the inherent complexity of clinical texts and may disrupt the learning process by introducing suboptimal sample ordering.
@@ V-C Analysis on Data Order
This trend indicates that starting with simpler samples helps models build a solid foundation, enabling them to better handle more complex cases later in training. In contrast, reversing the order disrupts this progression, leading to suboptimal performance. These findings highlight the importance of a well-structured curriculum that aligns with the natural progression of model learning.
@@ V-D Analysis on Clustering Number
The choice of clustering number directly impacts model performance and the interpretability of data distribution. As shown in the left panels of Figure 4, performance metrics across all three models and tasks achieve their peak when the clustering number is 3. This indicates that 3 clusters effectively balance the separation of data complexity levels, enabling models to learn optimally from progressively challenging samples.
The right panels of Figure 4 visualize the data distribution under different clustering numbers. When the clustering number is 3, the clusters are well-separated and intuitively aligned with the inherent structure of the data. Increasing the number of clusters beyond three results in over-segmentation, leading to redundant or overly granular clusters, while fewer clusters fail to capture sufficient complexity. These observations highlight the importance of selecting an optimal clustering number to achieve both strong performance and meaningful data partitioning.

### v2

@@ Abstract
Electronic medical records (EMRs) are crucial for modern healthcare, containing rich information about patient care, diagnoses, and treatments. However, their unstructured nature, domain-specific language, and complexity pose significant challenges for automated understanding. Existing methods often treat all data equally, limiting their ability to handle rare or complex cases effectively. We present TACL (Threshold-Adaptive Curriculum Learning), a novel framework that dynamically adjusts the training process based on sample complexity. Inspired by progressive learning, TACL categorizes data into difficulty levels, focusing on simpler cases early in training and gradually addressing more complex ones. A domain-specific pre-trained language model is used for difficulty assessment, considering semantic, syntactic, and contextual features. Additionally, TACL employs an adaptive training strategy to enhance task-specific performance and ensure generalization across diverse datasets. Experimental results on multilingual datasets, including MIMIC-III, MIMIC-IV, and Chinese clinical records, demonstrate TACL’s effectiveness in tasks such as ICD coding, readmission prediction, and TCM syndrome differentiation. TACL improves performance on rare and complex cases, providing a scalable and robust solution for medical text understanding.
@@ I Introduction
Medical texts, such as electronic medical records (EMRs), are essential for modern healthcare, capturing patient care, diagnoses, and treatments [giuffre2023harnessing]. However, their unstructured, domain-specific nature and variability—from straightforward cases to complex ones involving rare conditions—make automated understanding challenging.
As shown in Figure 1(a), EMRs exhibit multi-dimensional complexity, including narrative structure, clinical variability, and patient demographics. Traditional NLP models follow a “one-size-fits-all” approach, often failing on rare and complex cases [edin2023automated, gomes2024accurate, lu2023towards, ren2022hicu, vu2021label, zhao2024automated]. This motivates a key question: can models systematically learn from easier to harder cases, akin to human learning?
Curriculum learning [bengio2009curriculum], a training paradigm that progresses from simpler to more complex samples, has shown promise in healthcare applications like medical image segmentation [wang2023grenet, wang2023curriculum] and ICD coding [ren2022hicu]. However, key challenges remain: 1) How to effectively assess the difficulty of medical texts? and 2) How to develop adaptive strategies to generalize across diverse tasks and datasets? Simple metrics like sequence length or code frequency may fail to capture semantic nuances, as illustrated in Figure 1(b), necessitating more robust methods.
@@
To address these challenges, we propose TACL (Threshold-Adaptive Curriculum Learning), a framework designed to enhance medical text understanding. TACL introduces a novel difficulty assessment method leveraging a domain-specific pre-trained language model to quantify text complexity based on semantic, syntactic, and contextual features. Furthermore, it dynamically adjusts the curriculum to optimize learning across tasks and datasets.
@@
TACL is validated on multilingual datasets, including English (MIMIC-III [johnson2016mimic], MIMIC-IV [johnson2020mimic]) and Chinese clinical records, achieving state-of-the-art results in tasks like readmission prediction, out-hospital mortality, ICD coding, and TCM syndrome differentiation [ren2022tcm]. TACL excels in rare and complex cases, advancing medical text understanding.
Contributions
@@
  • First application of curriculum learning in medical text understanding: TACL introduces curriculum learning to handle the complexity and variability of clinical narratives, offering a novel perspective for this domain.
  • Difficulty assessment with domain-specific models: A pre-trained language model quantifies medical text complexity, enabling systematic difficulty categorization for curriculum learning.
@@
  • Adaptive training strategy: TACL dynamically adjusts the learning process, enhancing generalization across diverse medical datasets and improving robustness in real-world applications.
@@ II Related Work
Advancements in NLP have addressed challenges in automated medical text understanding, particularly in curriculum learning and medical text understanding.
@@ II-A Curriculum Learning in NLP
Curriculum learning [bengio2009curriculum] structures training by starting with simpler examples and progressing to harder ones. Dynamic methods [platanios-etal-2019-competence, huang2020curricularface, kong2021adaptive] adapt difficulty during training but focus on general NLP tasks, overlooking domain-specific challenges like medical texts. TACL bridges this gap with a threshold-adaptive mechanism tailored to medical data, dynamically prioritizing samples to enhance task performance and generalization.
@@ II-B Medical Text Understanding
Early rule-based approaches and traditional machine learning models required extensive feature engineering. Deep learning, including CNNs, RNNs, and transformer-based models like BERT [devlin2019bert], ClinicalBERT [huang2019clinicalbert], and BioBERT [lee2020biobert], improved medical NLP by leveraging pretraining on biomedical corpora. However, these models treat all samples equally, limiting their effectiveness on rare or complex cases. TACL addresses this by integrating curriculum learning to progressively train on simpler to harder cases, improving performance and generalization.
Datasets like MIMIC-III [johnson2016mimic] and MIMIC-IV [johnson2020mimic] highlight the variability in medical data, demonstrating the need for adaptive frameworks like TACL to handle heterogeneous and complex distributions.
@@ III-A Contextual Representation
  • English Tasks: For tasks such as ICD coding out-of-hospital mortality, and readmission prediction, we employ RoBERTa [lewis-etal-2020-pretrained] pre-trained on medical corpora.
  • Chinese Tasks: For TCM syndrome differentiation, we utilize ZY-BERT [ren2022tcm] pre-trained on traditional Chinese medicine datasets.
@@ IV-A Datasets
We evaluate TACL on four tasks using benchmark datasets partitioned into three difficulty levels (Easy, Medium, Hard). Table I summarizes dataset statistics.
Task 1: Patient Readmission: Predicts if a patient is readmitted within a specific period using MIMIC-III [johnson2016mimic]. Two subsets are used:
  • Discharge Summary: Detailed discharge notes.
  • Early Notes: Includes patient notes from 3 days earlier.
Task 2: Out-of-hospital Mortality: A binary classification task using MIMIC-IV [johnson2020mimic] discharge summaries to predict 30-day mortality post-discharge, excluding in-hospital deaths.
Task 3: ICD Coding: Multi-label classification on:
  • MIMIC-III Full (ICD-9): 52,723 documents, 8,929 codes (median: 14 codes/doc).
  • MIMIC-III 50 (ICD-9): 11,368 documents, top 50 codes (median: 5 codes/doc).
  • MIMIC-IV (ICD-10): Clinical notes annotated with granular ICD-10 codes.
Preprocessing includes lowercasing, removing non-alphanumeric characters, and truncation to 2,500–4,000 tokens.
Task 4: TCM Syndrome Differentiation: Classifies Traditional Chinese Medicine syndromes using TCM-SD [ren2022tcm], a Chinese-language dataset with domain-specific terminology.
@@ IV-B Implementation
Experiments are conducted on an NVIDIA L40 48GB GPU with fixed random seeds. TACL thresholds range from 0.5 to 0.8, optimized on validation subsets. Metrics include F1-score, AUC, and precision at K (P@K), computed on test sets using the best model on validation. All code will be publicly available [^1] https://anonymous.4open.science/r/medical-codes-354C.
@@ IV-C Baseline
We compare the proposed method against a range of strong baselines across all tasks. For tasks involving clinical and biomedical text, such as readmission prediction, out-of-hospital mortality, and ICD coding, we evaluate models like Bi-LSTM [schuster1997bidirectional], which utilizes pre-trained Word2Vec embeddings and bidirectional LSTM layers, and transformer-based models like BioBERT [alsentzer2019publicly] and ClinicalBERT [huang2019clinicalbert], both pre-trained on large biomedical corpora. For ICD coding, additional specialized models like CAML [mullenbach-etal-2018-explainable] with label attention mechanisms and MultiResCNN [li2020icd] for multi-scale feature extraction are employed, alongside PLM-ICD [huang-etal-2022-plm], which integrates a pre-trained BERT encoder with a label attention mechanism. For TCM syndrome differentiation, we evaluate Chinese-language models like RoBERTa-wwm [cui2021pre], pre-trained with Whole Word Masking on large-scale Chinese text, and ZY-BERT [ren2022tcm], a TCM-specific model trained on a large unlabelled corpus, achieving state-of-the-art performance on TCM tasks.
@@ V-A Main Results
Figure 3 summarizes TACL’s performance across all tasks, showing consistent improvements over traditional training methods in metrics such as Macro-F1, AUROC, and RP80.
TACL significantly enhances simpler models like Bi-GRU, CNN, and Bi-LSTM, particularly for readmission prediction and ICD coding, improving Macro-F1 scores and recall on rare labels. For advanced models like PLM-ICD, MultiResCNN, ClinicalBERT, and ZY-BERT, TACL provides smaller but consistent gains, especially for rare labels and imbalanced datasets, such as ICD-10 coding and TCM syndrome differentiation.
These results highlight TACL’s adaptability and scalability across models and tasks, demonstrating its potential to improve multilingual and domain-specific medical text understanding.
TABLE II: Performance on Patient Readmission Prediction.
@@ V-D Analysis on Clustering Number
Clustering number impacts performance and data interpretability. Metrics peak at 3 clusters (Figure 4), which balance data complexity and avoid over- or under-segmentation. Three clusters align well with inherent data structure, optimizing learning and performance.